OpenStudy (maheshmeghwal9):

Find: - \[\LARGE{\lim_{x \rightarrow 1}\frac{x^m-1}{x^n-1}}=?\] m & n are integers.

OpenStudy (maheshmeghwal9):

Answer is m/n. But how to get it ? Please help:)

OpenStudy (maheshmeghwal9):

@UnkleRhaukus :)

OpenStudy (unklerhaukus):

l'Hôpital's rule!

OpenStudy (anonymous):

well use this important identity\[x^k-1=(x-1)(x^{k-1}+x^{k-2}+...+x+1)\]

mathslover (mathslover):

\[\large{x^m -1 = (x-1)(x^{m-1}+x^{m-2}+..+x+1)}\] \[\large{x^n-1=(x-1)(x^{n-1}+x^{n-2}+...+x+1)}\] divide them

mathslover (mathslover):

x-1 will cancel out... and then there will be work of LIMITS..

OpenStudy (anonymous):

Let f(x)=x^m-1 then find f'(x) Let g(x)=x^n-1 then fing g'(x)

mathslover (mathslover):

\[\large{lim_{x \rightarrow 1}\frac{x^{m-1}+x^{m-2}+..+x+1}{x^{n-1}+x^{n-2}+..+x+1}}\] \[\large{\frac{1+1+..+1+1}{1+1+...1+1}}\] \[\large{\frac{m}{n}}\]

mathslover (mathslover):

in the numerator it will be : m times 1 and in the denominator it will be : n times 1 hence we have m/n as the answer

OpenStudy (anonymous):

Now, f'(x)=mx^(m-1) and g'(x)=nx^(n-1) Now,